# Graph Limits via Quotients

### v1

@@ 1.1.4 Equipartition-Consistent Random Graph Models (Section 5)
As our final contribution, we observe that random quotients $(\mathsf{G}_{k}[\mathsf{M}])_{k}$ of any grapheur $\mathsf{M}$ are related to each other by equipartitions of their vertex sets. Specifically, we observe that $\rho(d_{k,nk})\mathsf{G}_{nk}[\mathsf{M}]\overset{d}{=}\mathsf{G}_{k}[\mathsf{M}]$ for any equipartition $d_{k,nk}\colon[nk]\to[k]$, i.e., a map with equal-sized fibers. We call such sequences of random graphs equipartition-consistent random graph models, which play a role in our framework analogous to that played by consistent random graph models in graphon theory [lovasz2012random].
@@
Recall that grapheurs were defined as extremal random exchangeable measures in Definition 1.4. By Choquet’s theorem, any random exchangeable measure is therefore a mixture of grapheurs. Finally, we show in Proposition 5.1 that the sequence $(\mathsf{G}_{k}[\mathsf{M}])$ of quotients of a grapheur $\mathsf{M}$ converge back to $\mathsf{M}$ in a certain sense.
@@ 1.2 Related Work
The earliest two types of graph limits pertain to sequences of simple graphs that are dense, whose limits are given by graphons [LOVASZ2006933, convergent_seqs1], and degree-bounded, whose limits are given by graphings [benjamini_schramm, aldous2004objective, elek2007note]. We refer the reader to [lovasz2012large] for a survey of both limits. Both of these notions of graph limits are incomparable to ours, as both dense and degree-bounded sequences of simple graphs converge to the same limit in our framework, see Corollary 3.7. Further, dense graph limits can be characterized in terms of quotients [lovasz2012large, §12], called “right convergence” of graphons. However, these quotients are differently normalized compared to ours, and it is the set of quotients that converges in the graphon framework rather than their distribution as in our framework. Despite these distinctions, we highlight parallels between our development and the theory of graphons throughout our paper; in particular, grapheurs can be viewed as dual to graphons in a precise sense, see Section 3.3.
@@
A class of random graph models closely related to the ones we obtain by sampling edges from a grapheur is called edge-exchangeable [crane2016edge, Crane_Dempsey_2019, NIPS2016_1a0a283b, exchangeable_traits, janson2018edge]. These random graphs are obtained by adding random edges to a fixed infinite vertex set (often with repetition, yielding multigraphs), and can yield both sparse and dense random graphs [janson2018edge]. As explained in [janson2010graphons, Rmk. 4.4], every edge-exchangeable multigraph model can be obtained by sampling edges from a random exchangeable measure on $0,1]^{2}$. In fact, the random graphs we obtain by sampling edges in Section 4 are precisely these edge-exchangeable multigraphs, normalized to have unit edge weight, see Remark 4.2. From this perspective, convergence of grapheurs can be interpreted as convergence of the edge-exchangeable multigraph models associated to them, an interpretation we make precise in Proposition 4.1. Nevertheless, to our knowledge this notion of convergence of edge-exchangeable models has not been previously studied in the literature. In particular, the only metric between edge-exchangeable models we are aware of is an $\ell_{1}$ distance between edge probabilities [Crane_Dempsey_2019, Eq. (8)]. Convergence in this metric is too restrictive for our purposes, see Proposition 3.8. Further, our development of convergence via random quotients, which are not edge-exchangeable, allows us to take limits of general weighted graphs, to relate it to convergence of homomorphism densities and other graph parameters, and to test their values.
@@
The closest line of work to ours is the study of graphex processes and their associated graph limits and random graph models [caron_fox, veitch2015class, jmlr_graphon_proc, borgs2020identifiability, borgs2019sampling, multigraphex, JANSON2022103549]. Graphexes form a subset of random exchangeable measures on $\mathbb{R}^{2}_{\geq 0}$, and admit a characterization due to Kallenberg similarly to our grapheurs [kallenberg1990exchangeable]. Graphexes give rise to random graph models obtained by sampling edges from them [veitch2015class, borgs2019sampling, veitch2019sampling], analogously to our edge sampling procedure in Section 4.1. There are also several notions of limits for simple graphs and multigraphs, whereby such graphs converge to a limiting graphex [jmlr_graphon_proc, multigraphex]. These notions of limits are conjectured to be inequivalent [JANSON2022103549], and the space of graphexes with either topology is not compact. Some of these limits can be characterized by sampling random subgraphs (whose number of vertices is also random) [borgs2019sampling], while others can be characterized by convergence of differently-normalized homomorphism numbers [jmlr_graphon_proc, §2.5] (the authors of [jmlr_graphon_proc] normalize $\mathrm{hom}(H;G)$ by $(2|E(G)|)^{|V(H)|/2}$, whereas we normalize by $(2|E(G)|)^{|E(H)|}$). Graphex-based graph limits are again incomparable to ours. On the one hand, for a sequence of simple graphs our grapheur-based limits can only distinguish stars from a uniform component (Corollary 3.7), whereas graphex-based limits can capture more general limiting structures (see [JANSON2022103549, §9] or [multigraphex]). On the other hand, grapheur-based graph limits are defined for general sequences of nonnegatively-weighted graphs, for which graphex-based limits have not been defined, allowing us to test properties pertaining to the distribution of edge weight in the graph (see Section 4.3). In fact, grapheurs are particularly well-suited for property testing, since the space of grapheurs we study here is compact, in contrast to the space of graphexes; to our knowledge, there is no characterization of properties that can be tested using graphex-based limits. Another consequence of the compactness of the space of grapheurs is that our notion of graph limits can be characterized in several equivalent ways: by convergence of random quotients, convergence of random edge-sampled graphs, and convergence of homomorphism numbers and related graph parameters.
@@ 1.3 Notation and Preliminaries
where the infimum is over couplings $(X_{1}^{\prime},X_{2}^{\prime})$ of $X_{1}$ and $X_{2}$, the norm $\|\cdot\|_{1}$ is the entrywise $\ell_{1}$ norm, and the supremum is over 1-Lipschitz functions with respect to this norm. Note that a sequence of random graphs $(X_{n})\subseteq\Delta^{k\times k}$ converges weakly to the random graph $X\in\Delta^{k\times k}$ if and only if $\lim_{n}W_{1}(X_{n},X)=0$ by [villani2008optimal, Thm. 6.9].
For a compact Polish space $S$, the space $\mathcal{P}(S)$ endowed with the weak topology is itself compact and Polish. If $\mathsf{M}$ is a random measure, that is, a random variable taking values in the Polish space $\mathcal{P}(S)$, we write $\mathbb{E}_{\mathsf{M}}f$ for the scalar random variable obtained by integrating a deterministic function $f$ against the random measure $\mathsf{M}$. In particular, we write $\mathsf{M}(B)=\mathbb{E}_{\mathsf{M}}\mathbbm{1}_{B}$ to denote the scalar random variable giving the (random) measure of a measurable set $B$, where $\mathbbm{1}_{B}(x)=1$ if $x\in B$ and zero otherwise. A sequence of random measures $(\mathsf{M}_{n})$ converges weakly to $\mathsf{M}$ if $\mathrm{Law}(\mathsf{M}_{n})\to\mathrm{Law}(\mathsf{M})$ weakly, where these laws are probability measures over the space of probability measures. Equivalently, the sequence $(\mathsf{M}_{n})$ converges weakly to $\mathsf{M}$ if the sequence of scalar random variables $\mathbb{E}_{\mathsf{M}_{n}}f$ converges weakly to $\mathbb{E}_{\mathsf{M}}f$ for all continuous functions $f$ [kallenberg2005probabilistic, Thm. A2.3].
@@ 2 Quotient Convergence: Global and Local Views
This proves the first part of Theorem 1.2. To show the second part relating quotient convergence with convergence of homomorphism numbers, we express homomorphism numbers in terms of quotient densities and vice-versa. To this end, we begin by expressing quotient densities in terms of injective homomorphism numbers using [levin2025deFin, Prop. 6.8]. This expression is stated in terms of a partial order on multigraphs induced by quotients. Specifically, for two multigraphs $K\in\mathbb{N}^{n\times n},H\in\mathbb{N}^{m\times m}$ we say that $K$ is a refinement of $H$, denoted $K\leq_{R}H$, if $H$ is a quotient of $K$, i.e., if there exists a surjective map $f\colon[n]\to[m]$ such that $H=\rho(f)K$. We further let $R_{K,H}$ be the number of such maps $f$. Note that $R_{H,H}=|\mathrm{Aut}(H)|$ for any multigraph $H$. The poset of all multigraphs with the refinement order is a disjoint union of posets, one for each number of edges. The poset of all multigraphs with $n$ edges has a minimal element, namely, the disjoint union of $n$ edges, and a maximal element, namely, a single vertex with $n$ self-loops. We illustrate the poset of all multigraphs with two edges in Figure 3.
@@
The quotient density $t_{Q}(H;G)$ can be expressed in terms of a linear combination of injective homomorphism numbers over all refinements of $H$ [levin2025deFin, Prop. 6.8]:
@@
We are now ready to prove the second part of Theorem 1.2. The following argument is a simple modification of the one from [levin2025deFin, Prop. 6.8].
@@
Based on (11) and (10), we conclude that $(t_{Q}(H;G_{n}))$ converges for all multigraphs $H$ if and only if $(\mathrm{inj}(H;G_{n}))$ converges for all multigraphs $H$. Injective and non-injective homomorphism numbers can be similarly related, with one of the quantities being expressed in terms of linear combinations of the other as in [lovasz2012large, §5.2.3]. For example, we have [levin2025deFin, Eq. (73)] that
@@ 3.1 Characterizing Limits via Grapheurs
The second statement is a consequence of the compactness of $\mathcal{M}$, which follows from Proposition 3.3.
@@
We end this section by noting that valid polynomial inequalities in quotient densities precisely correspond to valid inequalities over grapheurs. We hope that grapheurs can be used to study such inequalities in future work, dually to the use of graphons in the study of inequalities in homomorphism densities [lovasz2012large, §16].
@@ 3.2 Structure of Grapheurs
Proposition 3.6 justifies the intuition outlined above: Up to a relabelling of the vertices of $G_{n}$, the entries of $E$ are entrywise limits of the edge weights of $G_{n}$, the entries of $\sigma,\varsigma$ are limits of the remaining degrees of vertices in $(G_{n})$ after removing edges with positive limiting weights, and the parameters $\vartheta,\theta$ capture the remaining edge weights in self loops and the rest of the edges, respectively. We prove Proposition 3.6 in Section 4.4 using an edge-sampling perspective on quotient convergence and a dual Szemerédi regularity lemma.
@@
Suppose $(G_{n}\in\mathbb{S}^{n})$ is a sequence of undirected graphs such that the sequence $(G_{n}/\mathbbm{1}^{\top}G_{n}\mathbbm{1})$ quotient-converges.
@@
    The limit is described by $(E,\sigma,\varsigma,\theta,\vartheta)$ with $\sigma=\varsigma$ and $E^{\top}=E$.
@@
    If $G_{n}\in\mathbb{S}^{n}_{\mathrm{sim}}$ is a sequence of simple graphs with $|E(G_{n})|\to\infty$, and is either dense or degree-bounded, then the sequence $(\frac{G_{n}}{2|E(G_{n})|})$ quotient converges to the grapheur $\lambda^{2}$.
@@
The first two claims follow directly from Proposition 3.6, since for any permutation $\pi_{n}\in\mathsf{S}_{\infty}$ the matrix $\pi_{n}G_{n}$ is, respectively, symmetric and has zero diagonal if $G_{n}$ is symmetric and has no self-loops. For the third claim, since $G_{n}$ is symmetric, we have $G_{n}^{\top}\mathbbm{1}=G_{n}\mathbbm{1}=\deg(G_{n})$ is the vector of degrees of the vertices in $G_{n}$. If $(G_{n})$ is dense, then $\frac{\max_{i}\deg(G_{n})_{i}}{2|E(G_{n})|}\leq\frac{c}{n}\xrightarrow{n\to\infty}0$ for some constant $c>0$. If $(G_{n})$ is degree-bounded, with all degrees bounded by $k$, then $\frac{\max_{i}\deg(G_{n})_{i}}{2|E(G_{n})|}\leq\frac{k}{2|E(G_{n})|}\xrightarrow{n\to\infty}0$. In either case, Proposition 3.6 shows that $\lim_{n}\frac{G_{n}}{2|E(G_{n})|}=\lambda^{2}$ since $\mathrm{diag}(G_{n})=0$ for all $n$. ∎
We end this section by giving a sufficient condition for the limit of a quotient-convergent sequence $(G_{n})$ to consist entirely of discrete hubs, i.e., such that $\sigma=\varsigma=0$ and $\theta=\vartheta=0$ in (5). In particular, this sufficient condition implies that the random quotients $(\rho(F_{k,n})G_{n})$ and the grapheurs $(\mathsf{M}_{G_{n}})$ converge not just weakly but also in total variation. Such total variation convergence was shown for graphs sampled both from graphons [lovasz2012large, Cor. 10.25] and from graphexes [borgs2019sampling, §5], and was the primary mode of convergence studied for edge-exchangeable random graph models [Crane_Dempsey_2019]. However, this condition is too restrictive for our purposes in this paper; for example, it does not apply to the sequence of stars in Example 2.2. In the following result, if $\mu,\nu$ are two discrete measures on a space $S$ then their total variation distance is $\mathrm{dist}_{\mathrm{TV}}(\mu,\nu)=\frac{1}{2}\sum_{x\in S}|\mu(\{x\})-\nu(\{x\})|$.

### v2

@@ 1.1.4 Equipartition-Consistent Random Graph Models (Section 5)
As our final contribution, we observe that random quotients $(\mathsf{G}_{k}[\mathsf{M}])_{k}$ of any grapheur $\mathsf{M}$ are related to each other by equipartitions of their vertex sets. Specifically, we observe that $\rho(d_{k,nk})\mathsf{G}_{nk}[\mathsf{M}]\overset{d}{=}\mathsf{G}_{k}[\mathsf{M}]$ for any equipartition $d_{k,nk}\colon[nk]\to[k]$, i.e., a map with equal-sized fibers. We call such sequences of random graphs equipartition-consistent random graph models, which play a role in our framework analogous to that played by consistent random graph models in graphon theory [33].
@@
Recall that grapheurs are extremal random exchangeable measures by Definition 1.4. By Choquet’s theorem, any random exchangeable measure is thus a mixture of grapheurs. Finally, we show in Proposition 5.1 that the sequence $(\mathsf{G}_{k}[\mathsf{M}])$ of quotients of a grapheur $\mathsf{M}$ converges back to $\mathsf{M}$ in a certain sense.
@@ 1.2 Related Work
The earliest two types of graph limits pertain to sequences of simple graphs that are dense, whose limits are given by graphons [32, 4], and degree-bounded, whose limits are given by graphings [3, 1, 18]. We refer the reader to [34] for a survey of both limits. Both of these notions of graph limits are incomparable to ours, as both dense and degree-bounded sequences of simple graphs converge to the same limit in our framework, see Corollary 3.7. Further, dense graph limits can be characterized in terms of quotients [34, §12], called “right convergence” of graphons. However, these quotients are differently normalized compared to ours, and it is the set of quotients that converges in the graphon framework rather than their distribution as in our framework. Despite these distinctions, we highlight parallels between our development and the theory of graphons throughout our paper; in particular, grapheurs can be viewed as dual to graphons in a precise sense, see Section 3.3.
@@
A class of random graph models closely related to the ones we obtain by sampling edges from a grapheur is called edge-exchangeable [15, 16, 10, 11, 24]. These random graphs are obtained by adding random edges to a fixed infinite vertex set (often with repetition, yielding multigraphs), and can yield both sparse and dense random graphs [24]. As explained in [23, Rmk. 4.4], every edge-exchangeable multigraph model can be obtained by sampling edges from a random exchangeable measure on $0,1]^{2}$. In fact, the random graphs we obtain by sampling edges in Section 4 are precisely these edge-exchangeable multigraphs, normalized to have unit edge weight, see Remark 4.2. From this perspective, convergence of grapheurs can be interpreted as convergence of the edge-exchangeable multigraph models associated to them, an interpretation we make precise in Proposition 4.1. Nevertheless, to our knowledge this notion of convergence of edge-exchangeable models has not been previously studied in the literature. In particular, the only metric between edge-exchangeable models we are aware of is an $\ell_{1}$ distance between edge probabilities [16, Eq. (8)]. Convergence in this metric is too restrictive for our purposes, see Proposition 3.8. Further, our development of convergence via random quotients, which are not edge-exchangeable, allows us to take limits of general weighted graphs, to relate it to convergence of homomorphism densities and other graph parameters, and to test their values.
@@
The closest line of work to ours is the study of graphex processes and their associated graph limits and random graph models [12, 45, 6, 5, 7, 8, 25]. Graphexes form a subset of random exchangeable measures on $\mathbb{R}^{2}_{\geq 0}$, and admit a characterization due to Kallenberg similarly to our grapheurs [28]. Graphexes give rise to random graph models obtained by sampling edges from them [45, 7, 44], analogously to our edge sampling procedure in Section 4.1. There are also several notions of limits for simple graphs and multigraphs, whereby such graphs converge to a limiting graphex [6, 8]. These notions of limits are conjectured to be inequivalent [25], and the space of graphexes with either topology is not compact. Some of these limits can be characterized by sampling random subgraphs (whose number of vertices is also random) [7], while others can be characterized by convergence of differently-normalized homomorphism numbers [6, §2.5] (the authors of [6] normalize $\mathrm{hom}(H;G)$ by $(2|E(G)|)^{|V(H)|/2}$, whereas we normalize by $(2|E(G)|)^{|E(H)|}$). Graphex-based graph limits are again incomparable to ours. On the one hand, for a sequence of simple graphs our grapheur-based limits can only distinguish stars from a uniform component (Corollary 3.7), whereas graphex-based limits can capture more general limiting structures (see [25, §9] or [8]). On the other hand, grapheur-based graph limits are defined for general sequences of nonnegatively-weighted graphs, for which graphex-based limits have not been defined, allowing us to test properties pertaining to the distribution of edge weight in the graph (see Section 4.3). In fact, grapheurs are particularly well-suited for property testing, since the space of grapheurs we study here is compact, in contrast to the space of graphexes; to our knowledge, there is no characterization of properties that can be tested using graphex-based limits. Another consequence of the compactness of the space of grapheurs is that our notion of graph limits can be characterized in several equivalent ways: by convergence of random quotients, convergence of random edge-sampled graphs, and convergence of homomorphism numbers and related graph parameters.
@@ 1.3 Notation and Preliminaries
where the infimum is over couplings $(X_{1}^{\prime},X_{2}^{\prime})$ of $X_{1}$ and $X_{2}$, the norm $\|\cdot\|_{1}$ is the entrywise $\ell_{1}$ norm, and the supremum is over 1-Lipschitz functions with respect to this norm. Note that a sequence of random graphs $(X_{n})\subseteq\Delta^{k\times k}$ converges weakly to the random graph $X\in\Delta^{k\times k}$ if and only if $\lim_{n}W_{1}(X_{n},X)=0$ by [46, Thm. 6.9].
For a compact Polish space $S$, the space $\mathcal{P}(S)$ endowed with the weak topology is itself compact and Polish. If $\mathsf{M}$ is a random measure, that is, a random variable taking values in the Polish space $\mathcal{P}(S)$, we write $\mathbb{E}_{\mathsf{M}}f$ for the scalar random variable obtained by integrating a deterministic function $f$ against the random measure $\mathsf{M}$. In particular, we write $\mathsf{M}(B)=\mathbb{E}_{\mathsf{M}}\mathbbm{1}_{B}$ to denote the scalar random variable giving the (random) measure of a measurable set $B$, where $\mathbbm{1}_{B}(x)=1$ if $x\in B$ and zero otherwise. A sequence of random measures $(\mathsf{M}_{n})$ converges weakly to $\mathsf{M}$ if $\mathrm{Law}(\mathsf{M}_{n})\to\mathrm{Law}(\mathsf{M})$ weakly, where these laws are probability measures over the space of probability measures. Equivalently, the sequence $(\mathsf{M}_{n})$ converges weakly to $\mathsf{M}$ if the sequence of scalar random variables $\mathbb{E}_{\mathsf{M}_{n}}f$ converges weakly to $\mathbb{E}_{\mathsf{M}}f$ for all continuous functions $f$ [26, Thm. A2.3].
@@ 2 Quotient Convergence: Global and Local Views
This proves the first part of Theorem 1.2. To show the second part relating quotient convergence with convergence of homomorphism numbers, we express homomorphism numbers in terms of quotient densities and vice-versa. To this end, we begin by expressing quotient densities in terms of injective homomorphism numbers using [31, Prop. 6.8]. This expression is stated in terms of a partial order on multigraphs induced by quotients. Specifically, for two multigraphs $K\in\mathbb{N}^{n\times n},H\in\mathbb{N}^{m\times m}$ we say that $K$ is a refinement of $H$, denoted $K\leq_{R}H$, if $H$ is a quotient of $K$, i.e., if there exists a surjective map $f\colon[n]\to[m]$ such that $H=\rho(f)K$. We further let $R_{K,H}$ be the number of such maps $f$. Note that $R_{H,H}=|\mathrm{Aut}(H)|$ for any multigraph $H$. The poset of all multigraphs with the refinement order is a disjoint union of posets, one for each number of edges. The poset of all multigraphs with $n$ edges has a minimal element, namely, the disjoint union of $n$ edges, and a maximal element, namely, a single vertex with $n$ self-loops. We illustrate the poset of all multigraphs with two edges in Figure 3.
@@
The quotient density $t_{Q}(H;G)$ can be expressed in terms of a linear combination of injective homomorphism numbers over all refinements of $H$ [31, Prop. 6.8]:
@@
We are now ready to prove the second part of Theorem 1.2. The following argument is a simple modification of the one from [31, Prop. 6.8].
@@
Based on (11) and (10), we conclude that $(t_{Q}(H;G_{n}))$ converges for all multigraphs $H$ if and only if $(\mathrm{inj}(H;G_{n}))$ converges for all multigraphs $H$. Injective and non-injective homomorphism numbers can be similarly related, with one of the quantities being expressed in terms of linear combinations of the other as in [34, §5.2.3]. For example, we have [31, Eq. (73)] that
@@ 3.1 Characterizing Limits via Grapheurs
The second statement follows from Proposition 3.3.
@@
We end this section by noting that valid polynomial inequalities in quotient densities precisely correspond to valid inequalities over grapheurs. We hope that grapheurs can be used to study such inequalities in future work, dually to the use of graphons in the study of inequalities in homomorphism densities [34, §16].
@@ 3.2 Structure of Grapheurs
Proposition 3.6 justifies the intuition outlined above: Up to a relabelling of the vertices of $G_{n}$, the entries of $E$ are entrywise limits of the edge weights of $G_{n}$, the entries of $\sigma,\varsigma$ are limits of the remaining degrees of vertices in $(G_{n})$ after removing edges with positive limiting weights, and the parameters $\vartheta,\theta$ capture the remaining edge weights in self loops and the rest of the edges, respectively. We prove Proposition 3.6 in Section 4.4 using an edge-sampling perspective on quotient convergence and an edge-based Szemerédi regularity lemma.
@@
Suppose $(G_{n}\in\mathbb{S}^{n}_{\geq 0})$ is a sequence of undirected graphs such that the sequence $(G_{n}/\mathbbm{1}^{\top}G_{n}\mathbbm{1})$ quotient-converges to a grapheur described by $(E,\sigma,\varsigma,\theta,\vartheta)$.
@@
    We have $\sigma=\varsigma$ and $E^{\top}=E$.
@@
    If $G_{n}\in\mathbb{S}^{n}_{\mathrm{sim}}$ is a sequence of simple graphs with $|E(G_{n})|\to\infty$, then $E=0$. If $(G_{n})$ is either dense or degree-bounded, then the sequence $(\frac{G_{n}}{2|E(G_{n})|})$ quotient converges to the grapheur $\lambda^{2}$.
@@
The first two claims follow directly from Proposition 3.6, since for any permutation $\pi_{n}\in\mathsf{S}_{\infty}$ the matrix $\pi_{n}G_{n}$ is, respectively, symmetric and has zero diagonal if $G_{n}$ is symmetric and has no self-loops. For the third claim, since each $G_{n}$ is simple, we have $G_{n}/\mathbbm{1}^{\top}G_{n}\mathbbm{1}\in\{0,1/n\}$ which converges entrywise to zero. Thus $E=0$ by Proposition 3.6. Since $G_{n}$ is symmetric, we have $G_{n}^{\top}\mathbbm{1}=G_{n}\mathbbm{1}=\deg(G_{n})$ is the vector of degrees of the vertices in $G_{n}$. If $(G_{n})$ is dense, then $\frac{\max_{i}\deg(G_{n})_{i}}{2|E(G_{n})|}\leq\frac{c}{n}\xrightarrow{n\to\infty}0$ for some constant $c>0$. If $(G_{n})$ is degree-bounded, with all degrees bounded by $k$, then $\frac{\max_{i}\deg(G_{n})_{i}}{2|E(G_{n})|}\leq\frac{k}{2|E(G_{n})|}\xrightarrow{n\to\infty}0$. In either case, Proposition 3.6 shows that $\lim_{n}\frac{G_{n}}{2|E(G_{n})|}=\lambda^{2}$ since $\mathrm{diag}(G_{n})=0$ for all $n$. ∎
We end this section by giving a sufficient condition for the limit of a quotient-convergent sequence $(G_{n})$ to consist entirely of discrete hubs, i.e., such that $\sigma=\varsigma=0$ and $\theta=\vartheta=0$ in (5). In particular, this sufficient condition implies that the random quotients $(\rho(F_{k,n})G_{n})$ and the grapheurs $(\mathsf{M}_{G_{n}})$ converge not just weakly but also in total variation. Such total variation convergence was shown for graphs sampled both from graphons [34, Cor. 10.25] and from graphexes [7, §5], and was the primary mode of convergence studied for edge-exchangeable random graph models [16]. However, this condition is too restrictive for our purposes in this paper; for example, it does not apply to the sequence of stars in Example 2.2. In the following result, if $\mu,\nu$ are two discrete measures on a space $S$ then their total variation distance is $\mathrm{dist}_{\mathrm{TV}}(\mu,\nu)=\frac{1}{2}\sum_{x\in S}|\mu(\{x\})-\nu(\{x\})|$.